# A Minimalist Optimizer Design for LLM Pretraining

### v2

@@ 5 Conclusion
Appendix A Appendix
A.1 Details of memory estimation for 1B and 7B models
Here we compute the memory estimate for both 1B and 7B LLaMA models. We only compute the major parameters, including embedding layers, attention and MLP layers. We follow prior works (zhao2024galore; han2024sltrain) in estimating the memory using bfloat16 format, where each floating point number occupies 2 bytes.
@@ A.1 Details of memory estimation for 1B and 7B models
A.2 Details of the experiments
For all LLaMA experiments, we follow zhao2024galore and set the sequence length to 256 and the batch size to 512, train using BF16 format and report evaluation perplexity as our metric. We also use a cosine learning rate with linear warm-up for the first 10% of the iterations. For low-rank optimizers (GaLore, Fira and APOLLO) we follow their suggested hyperparameters (including the rank) but tune the learning rates. For Muon we follow the implementation from liu2025muon.
We performed wandb sweeps for all methods we tested up to models of size 350M, searching learning rates within {0.00005, 0.0001, 0.0003, 0.0005, 0.001, 0.003, 0.005, 0.01}. For the 1B model, due to resource constraints we manually tune the learning rates using as starting point the optimal learning rate from the 350M sweep. For SCALE, we use learning rate 1e-3 for models sizes 60M, 130M and 350M, 2e-4 for the 1B and 1e-4 for the 7B model. In addition, our reported result for the 1B model uses the same learning rate scaling technique used by Muon (liu2025muon). Also, we set the last layer’s momentum parameter $\beta=0.9$, being a common choice for first order momentum. In addition, for all vector parameters we employ the Adam optimizer, following jordan2024muon; liao2024galore. This does not influence the memory usage because the vector parameters are orders of magnitude smaller in size compared to the matrix parameters.
A.3 Training Throughput Comparison of Different Methods
We conduct a throughput (tokens/sec) comparison of the different optimizers for training LLaMA 1B on 4xH100 GPUs using the same settings as specified in A.2. Table 7 shows that our method is about $18.5\%$ faster than singular-value (NS) based normalization methods (Muon/SWAN) and about $8\%$ faster than GaLore/Fira which project the Adam states to low-rank using SVD. Moreover, it achieves similar throughput to Adam/Stable-SPAM and APOLLO(-Mini).
@@ A.3 Training Throughput Comparison of Different Methods
Table 7: Throughput comparison of different methods for training LLaMA 1B on 4xH100 GPUs. $^{\star}:$ Due to code unavailability we cannot directly test the throughput of SWAN, instead we report the throughput using the Newton-Schulz (NS) approximation of Singular-value normalization (see jordan2024muon for details).
A.4 Pretraining Results on Additional LLM Architectures
For our main experiments we choose the LLaMA family of models as it is commonly used to benchmark memory efficient optimizers for LLM pretraining (such as in GaLore (zhao2024galore), Fira (chen2024fira), SWAN (ma2024swan) and APOLLO (zhu2025apollo)), which is the scope of our paper. However, to provide further evidence of the generality of our algorithm among LLM architectures, we conducted additional pretraining experiments with Qwen2-500M (yang2024qwen2technicalreport), GPT2-Medium (355M) (radford2019language) and Gemma-2B (team2024gemma). We followed the same experimental settings as in our LLaMA experiments, described in A.2. We provide the results bellow:
@@ A.4 Pretraining Results on Additional LLM Architectures
Tokens 40B
APOLLO 11.90 (9.09G)
SCALE 11.80 (6.06G)
@@
From table 9 we see that SCALE achieves a lower perplexity with an even smaller memory footprint than APOLLO, which further showcases its SOTA memory-efficient pretraining performance.
A.5 Results on Overtraining Regime
@@
Table 10: Results from training the 350M LLaMA model on C4 using different token budgets. Chinchilla ratio $1\times$ corresponds to (roughly) our default token budget for the given model, i.e., 7.8B tokens (following zhao2024galore), $2\times$ to 15.6B tokens and $4\times$ to 31.2B tokens. Our method maintains its SOTA memory-efficient pretraining performance among the different training budgets.
A.6 Fine-tuning Results
First, we want to emphasize that our optimizer design is focused on memory-efficient LLM pretraining. As pretraining is by far the most computationally expensive part of the LLM training process (for example pretraining LLaMA2-70B took about 1.72 million A100 GPU hours (touvron2023llama)), it is commonly believed that an efficient SOTA pretraining optimizer is of great importance.
Instead, for fine-tuning, existing Parameter-efficient fine-tuning (PEFT) methods already reduce the memory and compute requirements substantially while giving performance competitive to full fine-tuning. Therefore we don’t expect our optimizer to replace such approaches. However, to give further evidence of its generality, we provide some preliminary results on GLUE benchmarks with a pretrained RoBERTa-base (liu2019roberta) which we fine-tuned with our method and Adam (full fine-tuning), following the setup of zhao2024galore. We conduct a learning rate search for both methods and report the best epoch results (higher the better).
@@ A.6 Fine-tuning Results
Table 11: Results from fine-tuning RoBERTa-base on different GLUE benchmarks (wang2018glue). Memory consumption is indicated in the parenthesis.
@@
A.7 Learning Rate Sensitivity Analysis
@@
In Figure 5 we test the performance of our algorithm, SCALE, with different learning rates and compare it with that of Adam (Stable-SPAM version) (huang2025stablespam). We observe that both algorithms behave similarly with a reasonable range of learning rates.
A.8 Curves of Training Iteration versus Perplexity
@@
A.9 Discussion on why Column-wise matters for the Last Layer
@@
Figure 7: We present the column-norm values of the last layer’s gradient matrix at iterations 1,000 and 10,000 during the training of LLaMA 130M. The last layer of LLaMA 130M has gradient dimensions $d_{model}\times|V|=768\times 32,000$, which means that there are $|V|=32,000$ columns (and therefore $|V|$ corresponding column-norms). For clarity we limit the x axis to the first 1000 columns. We also note that due to the design of the SentencePiece tokenizer (kudo2018sentencepiece) used, lower token ids (and therefore lower last-layer column-ids) generally correspond to more frequent tokens (such as “_the”, “_and”, “_to”, etc., as we observed). Therefore, from the above figures it can be seen that more frequent tokens have much larger column-norms, potentially leading to imbalanced learning. We hypothesize that the effectiveness of column-wise is related to mitigating this phenomenon by normalizing all the column norms to the same level and thus enabling more balanced training dynamics.
@@
Table 12: Ablation results from training the 130M LLaMA model on C4 using “mixed” normalization schemes compared to SCALE.
@@
To explain the above results, our intuition suggests that the last layer in LLMs is somehow “special”. Indeed, it is the linear projection layer (often called Language modeling head) that maps the last hidden state (of dimension $d_{model}$) to the vocabulary logits (of dimension $|V|$ being the vocabulary size). Its gradient matrix shape is $d_{model}\times|V|$, i.e., it has $|V|$ columns, same as its weight matrix shape. Therefore, each gradient column of the last layer has a “physical meaning”, i.e., it corresponds to one of the $|V|$ vocabulary tokens.
We have observed that for certain (few) tokens, their corresponding last-layer column gradients have much larger norms compared to the rest. We have also observed that such columns correspond to more frequent tokens (see Figure 7 details). We hypothesize that such last-layer column-norm differences can potentially lead to uneven learning (i.e., more rare tokens receive smaller gradients and therefore are not “learned” by the model) or even training divergence (due to the large gradients corresponding to frequent tokens). In connection to this, kunstner2024heavy, aiming at explaining the large Adam-SGD gap in LLM training, observed that SGD is able to make much less progress on learning low-frequency classes compared to Adam. Based on those insights, we conclude that by normalizing the columns of the last layer our method obtains more stable training dynamics and a more “even” token learning behavior, even without using full per-parameter adaptivity as in Adam.
@@
A.10 Proofs for Section 2.2
In this section, we conduct the proof for Theorem 2.1. The proof follows liu2020improved. First, we have the following lemmas, which are variations of liu2020improved.
Lemma A.1.
Suppose that the assumptions in Theorem 2.1 hold. For SGD-M (2.7), we have
$\mathbb{E}\left[\left\|m_{l}^{t}-(1-\beta_{l})\sum_{i=1}^{t}\beta_{l}^{t-i}\nabla_{\theta_{l}}\ell(\theta^{i})\right\|^{2}\right]\leq\frac{1-\beta_{l}}{1+\beta_{l}}\left(1-\beta_{l}^{2t}\right)\sigma_{l}^{2}.$ (A.1)
Proof.  It is straightforward to see $m_{l}^{t}=(1-\beta_{l})\sum_{i=1}^{t}\beta_{l}^{t-i}g_{l}^{i}$, where $g_{l}^{i}:=\nabla_{\theta_{l}}f(\theta^{i};\xi_{i})$.
@@
Lemma A.2.
Suppose that the assumptions in Theorem 2.1 hold. For SGD-M (2.7), we have
$\mathbb{E}\left[\left\|\frac{1-\beta_{l}}{1-\beta_{l}^{t}}\sum_{i=1}^{t}\beta_{l}^{t-i}\nabla_{\theta_{l}}\ell(\theta^{i})-\nabla_{\theta_{l}}\ell(\theta^{t})\right\|^{2}\right]\leq\sum_{i=1}^{t-1}a_{l,t,i}\mathbb{E}\left[\left\|\theta^{i+1}-\theta^{i}\right\|^{2}\right],$ (A.2)
@@
$a_{l,t,i}=\frac{\gamma^{2}\beta_{l}^{t-i}}{1-\beta_{l}^{t}}\left(t-i+\frac{\beta_{l}}{1-\beta_{l}}\right).$ (A.3)
Proof.  Since
@@
then we get (A.2). ∎
Lemma A.3.
For SGD-M (2.7), define the auxiliary sequence by
$z_{l}^{t}:=\begin{cases}\theta_{l}^{t}&t=1\\ \frac{1}{1-\beta_{l}}\theta_{l}^{t}-\frac{\beta_{l}}{1-\beta_{l}}\theta_{l}^{t-1}&t\geq 2\end{cases}$ (A.4)
@@
Proof.  For $t=1$, we have (since $m^{0}=0$)
@@
Proposition A.1.
Suppose that the assumptions in Theorem 2.1 hold. For SGD-M (2.7), we have
$\mathbb{E}\left[\ell\left(z^{t+1}\right)\right]$ (A.5)
@@
where the auxiliary sequence $z_{l}^{t}$ is defined by (A.4).
Proof.  By Lemma A.3 we have that
@@
$\mathbb{E}_{\xi_{t}}\left[\ell\left(z^{t+1}\right)\right]\leq$ $\ell\left(z^{t}\right)+\frac{\rho_{0}}{2}\left(\sum_{l=1}^{L}\eta_{l}\right)\sum_{l=1}^{L}\left(\eta_{l}^{2}(\frac{\beta_{l}}{1-\beta_{l}})^{2}\gamma^{2}\mathbb{E}\|m_{l}^{t-1}\|^{2}\right)$ (A.6)
@@
Now from Lemma A.1 we have
@@
plug these into (A.6) we get:
@@
Proof. [Proof of Theorem 2.1] By Proposition A.1 and Lemma A.2, we have
$\mathbb{E}\left[\ell\left(z^{t+1}\right)\right]$ (A.7)
@@
$\phi^{t}=\ell\left(z^{t}\right)-\ell^{*}+\sum_{i=1}^{t-1}\sum_{l=1}^{L}c_{l,i}\left\|\theta_{l}^{t+1-i}-\theta_{l}^{t-i}\right\|^{2}$ (A.8)
where $c_{l,i}$ are constants to be determined.
From (A.7) we get
$\mathbb{E}\left[\phi^{t+1}\right]-\mathbb{E}\left[\phi^{t}\right]$ (A.9)
@@
$+\sum_{l=1}^{L}c_{l,1}\mathbb{E}\left\|\theta_{l}^{t+1}-\theta_{l}^{t}\right\|^{2}+\sum_{i=1}^{t-1}\sum_{l=1}^{L}(c_{l,i+1}-c_{l,i})\mathbb{E}\left[\left\|\theta_{l}^{t+1-i}-\theta_{l}^{t-i}\right\|^{2}\right]$
For the term $\sum_{l=1}^{L}c_{l,1}\mathbb{E}\left\|\theta_{l}^{t+1}-\theta_{l}^{t}\right\|^{2}$, we can bound it by (denote $\nabla_{l}^{t}=\nabla_{\theta_{l}}\ell(\theta^{t})$)
$\sum_{l=1}^{L}c_{l,1}\mathbb{E}\left\|\theta_{l}^{t+1}-\theta_{l}^{t}\right\|^{2}=\sum_{l=1}^{L}c_{l,1}\eta_{l}^{2}\mathbb{E}\left[\left\|m_{l}^{t}\right\|^{2}\right]$ (A.10)
$\leq$ $2\sum_{l=1}^{L}c_{l,1}\mathbb{E}\left[\left\|m_{l}^{t}-(1-\beta_{l})\sum_{i=1}^{t}c_{l,1}\beta_{l}^{t-i}\nabla_{l}^{i}\right\|^{2}\right]+2\sum_{l=1}^{L}\mathbb{E}\left[\left\|(1-\beta_{l})\sum_{i=1}^{t}\beta_{l}^{t-i}\nabla_{l}^{i}\right\|^{2}\right]$
$\leq$ $2\sum_{l=1}^{L}c_{l,1}\frac{1-\beta_{l}}{1+\beta_{l}}\sigma_{l}^{2}+2\sum_{l=1}^{L}c_{l,1}\mathbb{E}\left[\left\|(1-\beta_{l})\sum_{i=1}^{t}\beta_{l}^{t-i}\nabla_{l}^{i}\right\|^{2}\right]$
$\leq$ $\sum_{l=1}^{L}c_{l,1}\eta_{l}^{2}\left(2\frac{1-\beta_{l}}{1+\beta_{l}}\sigma_{l}^{2}+4\mathbb{E}\left[\left\|\nabla_{l}^{t}\right\|^{2}\right]\left(1-\beta_{l}^{t}\right)^{2}\right)+4\sum_{l=1}^{L}c_{l,1}(1-\beta_{l}^{t})^{2}\eta_{l}^{2}\mathbb{E}\left[\left\|\frac{1-\beta_{l}}{1-\beta_{l}^{t}}\sum_{i=1}^{t}\beta^{t-i}\nabla_{l}^{i}-\nabla_{l}^{t}\right\|^{2}\right]$
$\leq$ $\sum_{l=1}^{L}c_{l,1}\eta_{l}^{2}\left(2\frac{1-\beta_{l}}{1+\beta_{l}}\sigma_{l}^{2}+4\mathbb{E}\left[\left\|\nabla_{l}^{t}\right\|^{2}\right]\right)+4\sum_{l=1}^{L}c_{l,1}(1-\beta_{l}^{t})^{2}\eta_{l}^{2}\sum_{i=1}^{t-1}a_{l,t,i}\mathbb{E}\left[\left\|\theta^{i+1}-\theta^{i}\right\|^{2}\right]$
where for the last three inequalities we use (A.1) and

### v3

@@ 5 Conclusion
One limitation of our current study is that, due to computational constraints, we evaluate our optimizer on models up to 7B parameters and up to 100B training tokens. While such model sizes and token budgets are on par with comparisons in related literature in memory-efficient pretraining, an evaluation to even larger models and more tokens will be helpful to further show the benefits of the proposed optimizer.
Acknowledgements
M. Hong and A. Glentis are partly supported by NSF Awards ECCS-2426064 and EPCN-2311007.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. In particular, by making LLM pretraining more memory- and compute-efficient, the barrier to training larger models can be lowered. This has beneficial consequences including democratizing LLM development, but also includes potential risks from the wider access to powerful LLMs.
Appendix A LLM Acknowledgment
We acknowledge the use of LLMs for grammar checking and sentence polishing.
Appendix B Details of memory estimation for 1B and 7B models
Here we compute the memory estimate for both 1B and 7B LLaMA models. We only compute the major parameters, including embedding layers, attention and MLP layers. We follow prior works (Zhao et al., 2024; Han et al., 2024) in estimating the memory using bfloat16 format, where each floating point number occupies 2 bytes.
@@ Appendix B Details of memory estimation for 1B and 7B models
Appendix C Details of the experiments
For all LLaMA experiments, we follow (Zhao et al., 2024) and set the sequence length to 256 and the batch size to 512, train using BF16 format and report evaluation perplexity as our metric. We also use a cosine learning rate with linear warm-up for the first 10% of the iterations. For low-rank optimizers (GaLore, Fira and APOLLO) we follow their suggested hyperparameters (including the rank) but tune the learning rates. For Muon we follow the implementation from (Liu et al., 2025).
We performed wandb sweeps for all methods we tested up to models of size 350M, searching learning rates within {0.00005, 0.0001, 0.0003, 0.0005, 0.001, 0.003, 0.005, 0.01}. For the 1B model, due to resource constraints we manually tune the learning rates using as starting point the optimal learning rate from the 350M sweep. For SCALE, we use learning rate 1e-3 for models sizes 60M, 130M and 350M, 2e-4 for the 1B and 1e-4 for the 7B model. In addition, our reported result for the 1B model uses the same learning rate scaling technique used by Muon (Liu et al., 2025). Also, we set the last layer’s momentum parameter $\beta=0.9$, being a common choice for first-order momentum. In addition, for all vector parameters we employ the Adam optimizer, following (Jordan et al., 2024; Liao et al., 2024). This does not influence the memory usage because the vector parameters are orders of magnitude smaller in size compared to the matrix parameters.
Appendix D Training Throughput Comparison of Different Methods
We conduct a throughput (tokens/sec) comparison of the different optimizers for training LLaMA 1B on 4xH100 GPUs using the same settings as specified in C. Table 7 shows that our method is about $18.5\%$ faster than singular-value (NS) based normalization methods (Muon/SWAN) and about $8\%$ faster than GaLore/Fira which project the Adam states to low-rank using SVD. Moreover, it achieves similar throughput to Adam/Stable-SPAM and APOLLO(-Mini).
@@ Appendix D Training Throughput Comparison of Different Methods
Table 7: Throughput comparison of different methods for training LLaMA 1B on 4xH100 GPUs. $^{\star}:$ Due to code unavailability we cannot directly test the throughput of SWAN, instead we report the throughput using the Newton-Schulz (NS) approximation of Singular-value normalization (see (Jordan et al., 2024) for details).
Appendix E Ablation on Adding Momentum to the First Layer
We ablate using momentum both to the first and the last layer (SGD column-wise mmt-(first+last)) and include the results in the Table 8 bellow:
Method / LLaMA Model 60M 130M 350M
SGD column-wise (no momentum) 39.89 (0.12G) 28.85 (0.27G) 20.38 (0.74G)
SGD column-wise mmt-last (SCALE) 30.81 (0.15G) 22.57 (0.32G) 16.32 (0.80G)
SGD column-wise mmt-(first+last) 30.35 (0.18G) 22.26 (0.37G) 16.14 (0.87G)
Table 8: Experiments exploring the benefit of adding momentum to the embedding layer. The memory is included in parenthesis. The additional momentum for the first layer does not result to significant gains in model performance, which further validates our optimizer design choices.
Appendix F Pretraining Results on Additional LLM Architectures
For our main experiments we choose the LLaMA family of models as it is commonly used to benchmark memory efficient optimizers for LLM pretraining (such as in GaLore (Zhao et al., 2024), Fira (Chen et al., 2024), SWAN (Ma et al., 2025) and APOLLO (Zhu et al., 2025)), which is the scope of our paper. However, to provide further evidence of the generality of our algorithm among LLM architectures, we conducted additional pretraining experiments with Qwen2-500M (Yang et al., 2024), GPT2-Medium (355M) (Radford et al., 2019) and Gemma-2B (Team et al., 2024). We followed the same experimental settings as in our LLaMA experiments, described in C. We provide the results bellow:
@@ Appendix F Pretraining Results on Additional LLM Architectures
Tokens 32B
APOLLO 12.05 (9.09G)
SCALE 11.96 (6.06G)
@@
From table 10 we can see that our method achieves a lower perplexity with an even smaller memory footprint than APOLLO, which further showcases its SOTA performance as a memory-efficient pretraining algorithm.
Appendix G Results from Training LLaMA 7B up to 100B Tokens
Figure 5: Training 7B LLaMA on C4 up-to 100B tokens using the SCALE optimizer. The final evaluation perplexity achieved is 9.32.
To further validate the stability of SCALE in (near) industry-level training scope we pre-trained the LLaMA 7B model on C4 up-to 100B tokens (near the Chinchilla optimal training budget). We use the same settings as in our 7B experiment in Table 6, with the differences being scaling to two 8xH200 nodes with distributed data parallel, increased batch size of 1024 to fully utilize the GPU memory and learning rate increased by $\sqrt{2}$ to $\sqrt{2}\times 1e-4$ to account for the doubled batch size. Using the SCALE optimizer we were able to achieve a single-digit final evaluation perplexity of 9.32, while maintaining a stable loss trajectory throughout training (fully absent of loss spikes), as shown in Figure 5. This result validates our optimizer’s stability in large model and high budget runs.
Appendix H Results on Overtraining Regime
@@
Table 11: Results from training the 350M LLaMA model on C4 using different token budgets. Chinchilla ratio $1\times$ corresponds to (roughly) our default token budget for the given model, i.e., 7.8B tokens (following (Zhao et al., 2024)), $2\times$ to 15.6B tokens and $4\times$ to 31.2B tokens. Our method maintains its SOTA memory-efficient pretraining performance among the different training budgets.
Appendix I Fine-tuning Results
First, we want to emphasize that our optimizer design is focused on memory-efficient LLM pretraining. As pretraining is by far the most computationally expensive part of the LLM training process (for example pretraining LLaMA2-70B took about 1.72 million A100 GPU hours (Touvron et al., 2023)), it is commonly believed that an efficient SOTA pretraining optimizer is of great importance.
Instead, for fine-tuning, existing Parameter-efficient fine-tuning (PEFT) methods already reduce the memory and compute requirements substantially while giving performance competitive to full fine-tuning. Therefore we don’t expect our optimizer to replace such approaches. However, to give further evidence of its generality, we provide some preliminary results on GLUE benchmarks with a pretrained RoBERTa-base (Liu et al., 2019) which we fine-tuned with our method and Adam (full fine-tuning), following the setup of (Zhao et al., 2024). We conduct a learning rate search for both methods and report the best epoch results (higher the better).
@@ Appendix I Fine-tuning Results
Table 12: Results from fine-tuning RoBERTa-base on different GLUE benchmarks (Wang et al., 2018). Memory consumption is indicated in the parenthesis.
@@
Appendix J More Variance Analysis Results
In addition to our variance analysis in Figure 4, we ablate the impact of model size, architecture and tokenizer to the per-layer gradient variance behavior. As shown in Figures 6 and 7, The pattern remain similar to Figure 4, i.e., the last layer exhibits significantly more gradient variance than all other layers, and adding momentum to it helps reduce its variance.
Figure 6: Variance analysis experiment for a 350M LLaMA model on C4.
Figure 7: Variance analysis experiment for a GPT-125M model using the GPT-2 tokenizer on C4.
Appendix K Learning Rate Sensitivity Analysis
@@
In Figure 8 we test the performance of our algorithm, SCALE, with different learning rates and compare it with that of Adam (Stable-SPAM version) (Huang et al., 2025). We observe that both algorithms behave similarly with a reasonable range of learning rates.
Appendix L Curves of Training Iteration versus Perplexity
@@
Appendix M Discussion on why Column-wise matters for the Last Layer
@@
Figure 10: The column-norm values of the last layer’s gradient matrix at iterations 1,000 and 10,000 during the training of LLaMA 130M. The last layer of LLaMA 130M has gradient dimensions $d_{\textrm{model}}\times|V|=768\times 32,000$, which means that there are $|V|=32,000$ columns (and therefore $|V|$ corresponding column-norms). For clarity we limit the x axis to the first 1000 columns. We also note that due to the design of the SentencePiece tokenizer (Kudo and Richardson, 2018) used, lower token ids (and therefore lower last-layer column-ids) generally correspond to more frequent tokens (such as “_the”, “_and”, “_to”, etc., as we observed). Therefore, from the above figures it can be seen that more frequent tokens have much larger column-norms, potentially leading to imbalanced learning. We hypothesize that the effectiveness of column-wise is related to mitigating this phenomenon by normalizing all the column norms to the same level and thus enabling more balanced training dynamics.
@@
Table 13: Ablation results from training the 130M LLaMA model on C4 using “mixed” normalization schemes comapred to SCALE.
@@
To explain the above results, our intuition suggests that the last layer in LLMs is somehow “special”. Indeed, it is the linear projection layer (often called Language modeling head) that maps the last hidden state (of dimension $d_{\textrm{model}}$) to the vocabulary logits (of dimension $|V|$ being the vocabulary size). Its gradient matrix shape is $d_{model}\times|V|$, i.e., it has $|V|$ columns, same as its weight matrix shape. Therefore, each gradient column of the last layer has a “physical meaning”, i.e., it corresponds to one of the $|V|$ vocabulary tokens.
We have observed that for certain (few) tokens, their corresponding last-layer column gradients have much larger norms compared to the rest. We have also observed that such columns correspond to more frequent tokens (see Figure 10 details). We hypothesize that such last-layer column-norm differences can potentially lead to uneven learning (i.e., more rare tokens receive smaller gradients and therefore are not “learned” by the model) or even training divergence (due to the large gradients corresponding to frequent tokens). In connection to this, (Kunstner et al., 2024), aiming at explaining the large Adam-SGD gap in LLM training, observed that SGD is able to make much less progress on learning low-frequency classes compared to Adam. Based on those insights, we conclude that by normalizing the columns of the last layer our method obtains more stable training dynamics and a more “even” token learning behavior, even without using full per-parameter adaptivity as in Adam.
@@
Appendix N Proofs for Section 2.2
In this section, we conduct the proof for Theorem 2.1. The proof follows Liu et al. (2020). First, we have the following lemmas, which are variations of Liu et al. (2020, Lemma 1 and 2).
Lemma N.1.
Suppose that the assumptions in Theorem 2.1 hold. For SGD-M (7), we have
$\mathbb{E}\left[\left\|m_{l}^{t}-(1-\beta_{l})\sum_{i=1}^{t}\beta_{l}^{t-i}\nabla_{\theta_{l}}\ell(\theta^{i})\right\|^{2}\right]\leq\frac{1-\beta_{l}}{1+\beta_{l}}\left(1-\beta_{l}^{2t}\right)\sigma_{l}^{2}.$ (12)
Proof.
It is straightforward to see $m_{l}^{t}=(1-\beta_{l})\sum_{i=1}^{t}\beta_{l}^{t-i}g_{l}^{i}$, where $g_{l}^{i}:=\nabla_{\theta_{l}}f(\theta^{i};\xi_{i})$.
@@
Lemma N.2.
Suppose that the assumptions in Theorem 2.1 hold. For SGD-M (7), we have
$\mathbb{E}\left[\left\|\frac{1-\beta_{l}}{1-\beta_{l}^{t}}\sum_{i=1}^{t}\beta_{l}^{t-i}\nabla_{\theta_{l}}\ell(\theta^{i})-\nabla_{\theta_{l}}\ell(\theta^{t})\right\|^{2}\right]\leq\sum_{i=1}^{t-1}a_{l,t,i}\mathbb{E}\left[\left\|\theta^{i+1}-\theta^{i}\right\|^{2}\right],$ (13)
@@
$a_{l,t,i}=\frac{\gamma^{2}\beta_{l}^{t-i}}{1-\beta_{l}^{t}}\left(t-i+\frac{\beta_{l}}{1-\beta_{l}}\right).$ (14)
Proof.
Since
@@
then we get (13). ∎
Lemma N.3.
For SGD-M (7), define the auxiliary sequence by
$z_{l}^{t}:=\begin{cases}\theta_{l}^{t}&t=1\\ \frac{1}{1-\beta_{l}}\theta_{l}^{t}-\frac{\beta_{l}}{1-\beta_{l}}\theta_{l}^{t-1}&t\geq 2\end{cases}$ (15)
@@
Proof.
For $t=1$, we have (since $m^{0}=0$)
@@
Proposition N.4.
Suppose that the assumptions in Theorem 2.1 hold. For SGD-M (7), we have
$\mathbb{E}\left[\ell\left(z^{t+1}\right)\right]$ (16)
@@
where the auxiliary sequence $z_{l}^{t}$ is defined by (15).
Proof.
By Lemma N.3 we have that
@@
$\mathbb{E}_{\xi_{t}}\left[\ell\left(z^{t+1}\right)\right]\leq$ $\ell\left(z^{t}\right)+\frac{\rho_{0}}{2}\left(\sum_{l=1}^{L}\eta_{l}\right)\sum_{l=1}^{L}\left(\eta_{l}^{2}(\frac{\beta_{l}}{1-\beta_{l}})^{2}\gamma^{2}\mathbb{E}\|m_{l}^{t-1}\|^{2}\right)$ (17)
@@
Now from Lemma N.1 we have
@@
plug these into (17) we get:
@@
Proof of Theorem 2.1.
By Proposition N.4 and Lemma N.2, we have
$\mathbb{E}\left[\ell\left(z^{t+1}\right)\right]$ (18)
@@
$\phi^{t}=\ell\left(z^{t}\right)-\ell^{*}+\sum_{i=1}^{t-1}c_{i}\left\|\theta^{t+1-i}-\theta^{t-i}\right\|^{2}$ (19)
where $c_{i}$ are constants to be determined.
From (18) we get
$\mathbb{E}\left[\phi^{t+1}\right]-\mathbb{E}\left[\phi^{t}\right]$ (20)
@@
$+c_{1}\mathbb{E}\left\|\theta^{t+1}-\theta^{t}\right\|^{2}+\sum_{i=1}^{t-1}(c_{i+1}-c_{i})\mathbb{E}\left[\left\|\theta^{t+1-i}-\theta^{t-i}\right\|^{2}\right]$
For the term $c_{1}\mathbb{E}\left\|\theta^{t+1}-\theta^{t}\right\|^{2}$, we can bound it by (denote $\nabla_{l}^{t}=\nabla_{\theta_{l}}\ell(\theta^{t})$)
$c_{1}\mathbb{E}\left\|\theta^{t+1}-\theta^{t}\right\|^{2}=\sum_{l=1}^{L}c_{1}\mathbb{E}\left\|\theta_{l}^{t+1}-\theta_{l}^{t}\right\|^{2}=\sum_{l=1}^{L}c_{1}\eta_{l}^{2}\mathbb{E}\left[\left\|m_{l}^{t}\right\|^{2}\right]$ (21)
$\leq$ $2\sum_{l=1}^{L}c_{1}\mathbb{E}\left[\left\|m_{l}^{t}-(1-\beta_{l})\sum_{i=1}^{t}c_{1}\beta_{l}^{t-i}\nabla_{l}^{i}\right\|^{2}\right]+2\sum_{l=1}^{L}\mathbb{E}\left[\left\|(1-\beta_{l})\sum_{i=1}^{t}\beta_{l}^{t-i}\nabla_{l}^{i}\right\|^{2}\right]$
$\leq$ $2\sum_{l=1}^{L}c_{1}\frac{1-\beta_{l}}{1+\beta_{l}}\sigma_{l}^{2}+2\sum_{l=1}^{L}c_{1}\mathbb{E}\left[\left\|(1-\beta_{l})\sum_{i=1}^{t}\beta_{l}^{t-i}\nabla_{l}^{i}\right\|^{2}\right]$
$\leq$ $\sum_{l=1}^{L}c_{1}\eta_{l}^{2}\left(2\frac{1-\beta_{l}}{1+\beta_{l}}\sigma_{l}^{2}+4\mathbb{E}\left[\left\|\nabla_{l}^{t}\right\|^{2}\right]\left(1-\beta_{l}^{t}\right)^{2}\right)+4\sum_{l=1}^{L}c_{1}(1-\beta_{l}^{t})^{2}\eta_{l}^{2}\mathbb{E}\left[\left\|\frac{1-\beta_{l}}{1-\beta_{l}^{t}}\sum_{i=1}^{t}\beta^{t-i}\nabla_{l}^{i}-\nabla_{l}^{t}\right\|^{2}\right]$
$\leq$ $\sum_{l=1}^{L}c_{1}\eta_{l}^{2}\left(2\frac{1-\beta_{l}}{1+\beta_{l}}\sigma_{l}^{2}+4\mathbb{E}\left[\left\|\nabla_{l}^{t}\right\|^{2}\right]\right)+4\sum_{l=1}^{L}c_{1}(1-\beta_{l}^{t})^{2}\eta_{l}^{2}\sum_{i=1}^{t-1}a_{l,t,i}\mathbb{E}\left[\left\|\theta^{i+1}-\theta^{i}\right\|^{2}\right]$
where for the last three inequalities we use (12) and